# QRAM architectures using 3D superconducting cavities

### v2

@@ II Bucket-Brigade QRAM
Figure 1: Comparison of fanout and bucket-brigade QRAM with two-level routers. a) By constructing a tree of quantum routers, one can access data in superposition. We consider the state of the tree in a query to $D_{4}$ and $D_{5}$ (highlighted) where the indicated router experiences an amplitude-damping error. b) In a fanout approach [2], all routers are set at each level according to the state of the address at that level, leading to a GHZ-like state. The resulting state of the QRAM tree is thus highly entangled and susceptible to decoherence. An amplitude-damping event on the indicated router causes the state of the tree to collapse and the query to fail. c) In the case of bucket-brigade QRAM [15, 18], only routers participating in a query to a given branch are activated. This leads to a less entangled state of the QRAM tree as compared to fanout QRAM. Amplitude damping now can not occur on the router indicated in a) and the superposition state is not disrupted.
@@ III.1 Gate protocol
The additional single-cavity rotation in $\textsf{C}_{0}\textsf{SWAP}$ is to correct for unwanted phases. See Appendix B for an explicit verification that the gate functions as intended on the states of interest $|n_{a}00\rangle,|n_{a}01\rangle,|n_{a}10\rangle,n_{a}=0,1$, where the third index refers to occupation in mode $c$. We note that these CSWAP gates do not behave as expected if modes $b$ and $c$ are both initially occupied. We might expect that the overall operation should be trivial. However, due both to the Hong-Ou-Mandel effect [36] and to phase shifts due to the $\mathsf{CZ}_{ab}$ gate, population is transferred out of the computational subspace. We stress that in the absence of thermal photons, we never expect these modes to be simultaneously occupied in the course of a QRAM query. Thus all population should ideally remain in the computational subspace  [^1] We are assuming that the QRAM device is initialized in the vacuum state, as opposed to some proposed QRAM architectures that can be initialized in an arbitrary state [18]. Nevertheless, both modes may become occupied due to thermal excitations, thus we discuss this case in Appendix B.
@@
We emphasize that the error detection of ancilla errors is performed after every $\mathsf{CZ},\textsf{CSWAP}$ gate. No “which path” information is revealed by performing measurements on the ancillas, as they are disentangled from the tree after every gate and do not collapse the superposition of queries. This is to be contrasted with error detection of cavity photon-loss events in the dual-rail case, which must wait until the address and bus qubits have been routed out of the tree. Such measurements would reveal which path information and collapse the superposition state of the tree. For more details on error detection of the $\mathsf{CZ}$ operation, see Appendix C.
@@ III.3 Infidelity and error mitigation by post selection
Table 1: Parameters used in numerical simulations and associated jump operators. The decay and dephasing rates of the cavities are $\Gamma_{1}^{c}=1/T_{1}^{c}$, $\Gamma_{\phi}^{c}=1/T_{\phi}^{c}$, respectively, implicitly defining the associated coherence times. For the transmon, decay and dephasing rates are generally state dependent [44]. We make the usual assumption of bosonic enhancement these rates [24], taking $\Gamma_{\phi}^{\,t,gg}=0,\Gamma_{\phi}^{\,t,ff}=4\Gamma_{\phi}^{\,t,ee}$ and $\Gamma_{1}^{\,t,ef}=2\Gamma_{1}^{\,t,ge}$. The associated coherence times are defined as e.g. $T_{1}^{\,t,ef}=1/\Gamma_{1}^{\,t,ef}$.
Finally, we assume a thermal population of $n_{\mathrm{th}}=0.01$. jump op. PS1 [41] PS2 PS3 $T_{1}^{\,c}$ $\hat{a}$ 0.6 ms 25 ms [19] 25 ms $T_{\phi}^{\,c}$ $\hat{a}^{\dagger}\hat{a}$ 5 ms 106 ms [19] 106 ms $T_{1}^{\,t,ge}$ $|g\rangle\langle e|$ 0.2 ms 0.5 ms [43] 2 ms $T_{\phi}^{\,t,ee}$ $|e\rangle\langle e|$ 0.4 ms 0.9 ms [43] 4 ms
@@
where $t_{\mathsf{CZ}}N_{\rm ts}$ is the time per query and $1/P_{\mathrm{no\,flag}}$ is the expected number of trials until success. The success rate is then $\Gamma_{\rm success}=1/T_{\rm success}$ In the dual-rail case $P_{\mathrm{no\,flag}}$ includes the probability of a photon-loss event for the address or bus qubits during the course of a query.
@@ VI Acknowledgments
S.P. and S.M.G. receive consulting fees and are equity holders in Quantum Circuits, Inc.
@@ Appendix C Fidelity of the $\mathsf{CZ}$ operations
In the dual-rail case we can also detect photon-loss events in the cavities by performing a parity check [25]. It is important to note that this parity check can be done only at the end of the QRAM query, once the addresses and bus have been routed in and back out of the tree  [^4] Note that the QRAM query times are shorter than $T_{1}^{c}$ for the parameters considered in this paper. Thus a parity check performed only at the end of the circuit is sufficient, and we do not expect to be limited by uncaught cavity decay/heating events [71]. Otherwise, we reveal “which path” information which destroys the superposition state. Thus in the following we do not explicitly simulate a parity check at the completion of the $\mathsf{CZ}$ operation, which would take additional time $\sim\pi/\chi$ [25, 24] and unrealistically increase our error probability per time step due to additional uncaught transmon errors. Instead, we perform the ancilla measurement(s) as described above, then project onto the dual-rail basis to obtain an estimate of the fidelity boost due to utilizing dual-rail qubits. The overall no-error measurement operator is
@@ Appendix D GUE state transfer
$+\sum_{\begin{subarray}{c}\mu=b,c\\ j=1,2,3,4\end{subarray}}\omega_{\mu j}\hat{\mu}_{j}^{\dagger}\hat{\mu}_{j},$ (38)
@@
$\quad-\frac{i}{2}\sum_{\begin{subarray}{c}d=\mathrm{R,L}\\ \mu=b,c\end{subarray}}[\hat{L}_{d}^{(\mu)}(\omega_{0})]^{\dagger}\hat{L}_{d}^{(\mu)}(\omega_{0})$ (48)
@@
$\quad-\frac{i}{2}\sum_{\begin{subarray}{c}d=\mathrm{R,L}\\ \mu=a,b,c\end{subarray}}[\hat{L}_{d}^{(\mu)}(\omega_{0})]^{\dagger}\hat{L}_{d}^{(\mu)}(\omega_{0})$ (59)

### v3

# Quantum random access memory architectures using 3D superconducting cavities

@@ II Bucket-Brigade QRAM
Figure 1: Comparison of fanout and bucket-brigade QRAM with two-level routers. a) By constructing a tree of quantum routers, one can access data in superposition. We consider the state of the tree in a query to $D_{4}$ and $D_{5}$ (highlighted) where the indicated router experiences an amplitude-damping error (the general argument holds also for other error types, see Ref. [18]). b) In a fanout approach [2], all routers are set at each level according to the state of the address at that level, leading to a GHZ-like state. The resulting state of the QRAM tree is thus highly entangled and susceptible to decoherence. An amplitude-damping event on the indicated router causes the state of the tree to collapse and the query to fail. c) In the case of bucket-brigade QRAM [15, 18], only routers participating in a query to a given branch are activated. This leads to a less entangled state of the QRAM tree as compared to fanout QRAM. Amplitude damping now can not occur on the router indicated in a) and the superposition state is not disrupted.
@@ III.1 Gate protocol
The additional single-cavity rotation in $\textsf{C}_{0}\textsf{SWAP}$ is to correct for unwanted phases. See Appendix B for an explicit verification that the gate functions as intended on the states of interest $|n_{a}00\rangle,|n_{a}01\rangle,|n_{a}10\rangle,n_{a}=0,1$, where the third index refers to occupation in mode $c$. We note that these CSWAP gates do not behave as expected if modes $b$ and $c$ are both initially occupied. We might expect that the overall operation should be trivial. However, due both to the Hong-Ou-Mandel (HOM) effect [36] and to phase shifts due to the $\mathsf{CZ}_{ab}$ gate, population is transferred out of the computational subspace. We stress that in the absence of thermal photons, we never expect these modes to be simultaneously occupied in the course of a QRAM query. Thus all population should ideally remain in the computational subspace  [^1] We are assuming that the QRAM device is initialized in the vacuum state, as opposed to some proposed QRAM architectures that can be initialized in an arbitrary state [18]. Nevertheless, both modes may become occupied due to thermal excitations, thus we discuss this case in Appendix B.
@@
We emphasize that the error detection of ancilla errors is performed after every $\mathsf{CZ},\textsf{CSWAP}$ gate, see Appendix C. No “which path” information is revealed by performing measurements on the ancillas, as they are disentangled from the tree after every gate and do not collapse the superposition of queries. This is to be contrasted with error detection of cavity photon-loss events in the dual-rail case, which must wait until the address and bus qubits have been routed out of the tree. We cannot perform mid-circuit cavity-photon-loss detection because we initialize the tree in the vacuum state, as opposed to dual-rail logical states. Thus, such error detection would reveal which path information and collapse the superposition state of the tree. For example, detecting that a photon is present in a router (indicating occupation of a dual-rail logical state) destroys those terms in the superposition state of the tree where this router was in vacuum. If we hypothetically had access to a CSWAP operation that performs as expected when both target modes are occupied, then we may initialize the tree in logical dual-rail states and perform mid circuit error detection of cavity photon loss. Indeed, detecting cavity photon-loss errors becomes important in the large $n$ limit, when query times approach cavity lifetimes $T_{1}^{c}$. Thus, constructing such a CSWAP operation is an interesting avenue of further research.
@@ III.2 Resource estimates
The time to complete a QRAM query is then $t_{n}=t_{\mathsf{CZ}}N_{\rm ts}$. For the parameters considered in this work, $t_{n}$ is generally short when compared with $T_{1}^{c}$. For instance, for $n=12$ and $N=2^{12}$, we obtain $t_{12}=130\mu$s.
@@ III.3 Infidelity and error mitigation by post selection
Table 1: Parameters used in numerical simulations and associated jump operators. The decay and dephasing rates of the cavities are $\Gamma_{1}^{c}=1/T_{1}^{c}$, $\Gamma_{\phi}^{c}=1/T_{\phi}^{c}$, respectively, implicitly defining the associated coherence times. For the transmon, decay and dephasing rates are generally state dependent [44]. We make the usual assumption of bosonic enhancement these rates [24], taking $\Gamma_{\phi}^{\,t,gg}=0,\Gamma_{\phi}^{\,t,ff}=4\Gamma_{\phi}^{\,t,ee}$ and $\Gamma_{1}^{\,t,ef}=2\Gamma_{1}^{\,t,ge}$. The associated coherence times are defined as e.g. $T_{1}^{\,t,ef}=1/\Gamma_{1}^{\,t,ef}$. Finally, we assume a thermal population of $n_{\mathrm{th}}=0.01$.
jump op. PS1 [41] PS2 PS3
$T_{1}^{\,c}$ $\hat{a}$ 0.6 ms 25 ms [19] 25 ms
$T_{\phi}^{\,c}$ $\hat{a}^{\dagger}\hat{a}$ 5 ms 106 ms [19] 106 ms
$T_{1}^{\,t,ge}$ $|g\rangle\langle e|$ 0.2 ms 0.5 ms [43] 2 ms
$T_{\phi}^{\,t,ee}$ $|e\rangle\langle e|$ 0.4 ms 0.9 ms [43] 4 ms
@@
where $t_{\mathsf{CZ}}N_{\rm ts}$ is the time per query and $1/P_{\mathrm{no\,flag}}$ is the expected number of trials until success. The success rate is then $\Gamma_{\rm success}=1/T_{\rm success}$. In the dual-rail case $P_{\mathrm{no\,flag}}$ includes the probability of a photon-loss event for the address or bus qubits during the course of a query.
@@ VI Acknowledgments
S.P. and S.M.G. receive consulting fees from Quantum Circuits, Inc. S.M.G. is an equity holder in Quantum Circuits, Inc.
@@ Appendix C Fidelity of the $\mathsf{CZ}$ operations
In the dual-rail case we can also detect photon-loss events in the cavities by performing a parity check [25]. It is important to note that this parity check can be done only at the end of the QRAM query, once the addresses and bus have been routed in and back out of the tree  [^4] Note that the QRAM query times $t_{n}=t_{\mathsf{CZ}}N_{\rm ts}$ are shorter than $T_{1}^{c}$ for the parameters considered in this paper. Thus a parity check performed only at the end of the circuit is sufficient, and we do not expect to be limited by uncaught cavity decay/heating events [71]. Otherwise, we reveal “which path” information which destroys the superposition state. Thus in the following we do not explicitly simulate a parity check at the completion of the $\mathsf{CZ}$ operation, which would take additional time $\sim\pi/\chi$ [25, 24] and unrealistically increase our error probability per time step due to additional uncaught transmon errors. Instead, we perform the ancilla measurement(s) as described above, then project onto the dual-rail basis to obtain an estimate of the fidelity boost due to utilizing dual-rail qubits. The overall no-error measurement operator is
@@ Appendix D GUE state transfer
$+\sum_{\begin{subarray}{c}\mu=b,c\\ j=1,2,3,4\end{subarray}}\omega_{\mu j}\hat{\mu}_{j}^{\dagger}\hat{\mu}_{j},$
@@
$\quad-\frac{i}{2}\sum_{\begin{subarray}{c}d=\mathrm{R,L}\\ \mu=b,c\end{subarray}}[\hat{L}_{d}^{(\mu)}(\omega_{0})]^{\dagger}\hat{L}_{d}^{(\mu)}(\omega_{0})$
@@
$\quad-\frac{i}{2}\sum_{\begin{subarray}{c}d=\mathrm{R,L}\\ \mu=a,b,c\end{subarray}}[\hat{L}_{d}^{(\mu)}(\omega_{0})]^{\dagger}\hat{L}_{d}^{(\mu)}(\omega_{0})$